8.4

2.)

Construct a 90% confidence interval for (P1 – P2) in each of the following situations. a. ny = 400; 21 = 0.64; n2 = 400; P2 =90% Confidence interval for b.= ____,____

90% Confidence interval for c.= ____,____
Answer #1

Confidence interval =(\hat{p_{1}}-\hat{p_{2}})\pm z*\sqrt{\frac{\hat{p_{1}}*\hat{q_{1}}}{n_{1}}+\frac{\hat{p_{2}}*\hat{q_{2}}}{n_{2}}}

a) 11 = 400 , 2 = 400 , //img.homeworklib.com/questions/28e1c2a0-b288-11eb-882f-0923a640395e.png?x-oss-process=image/resize,w_560 1= 0.64, //img.homeworklib.com/questions/292024b0-b288-11eb-ae58-bfa2d3f54989.png?x-oss-process=image/resize,w_560 2 = 0.56 , //img.homeworklib.com/questions/295c52e0-b288-11eb-82c8-7d9cde331299.png?x-oss-process=image/resize,w_560 1 = 1 – //img.homeworklib.com/questions/29bb8e20-b288-11eb-9311-f7352a6b4900.png?x-oss-process=image/resize,w_560 1 = 0.36 , //img.homeworklib.com/questions/29f830c0-b288-11eb-953d-176cbb53a276.png?x-oss-process=image/resize,w_560 2 = 1 – //img.homeworklib.com/questions/2a350180-b288-11eb-8542-65af756c87c3.png?x-oss-process=image/resize,w_560 2 = 0.44

Confidence level = 0.90

α = 1 - 0.90 = 0.1 , 1 - (α/2) = 0.95

So z = 1.645 ----( from z score table )

=(0.64-0.56)\pm 1.645*\sqrt{\frac{(0.64*0.36)}{400}+\frac{(0.56*0.44)}{400}}

= 0.023 and 0.137

90% confidence interval for P1-P2 is 0.023 , 0.137

b) 11 = 180 , 2 = 250 , //img.homeworklib.com/questions/28e1c2a0-b288-11eb-882f-0923a640395e.png?x-oss-process=image/resize,w_560 1= 0.28, //img.homeworklib.com/questions/292024b0-b288-11eb-ae58-bfa2d3f54989.png?x-oss-process=image/resize,w_560 2 = 0.25 , //img.homeworklib.com/questions/295c52e0-b288-11eb-82c8-7d9cde331299.png?x-oss-process=image/resize,w_560 1 = 1 – //img.homeworklib.com/questions/29bb8e20-b288-11eb-9311-f7352a6b4900.png?x-oss-process=image/resize,w_560 1 = 0.72 , //img.homeworklib.com/questions/29f830c0-b288-11eb-953d-176cbb53a276.png?x-oss-process=image/resize,w_560 2 = 1 – //img.homeworklib.com/questions/2a350180-b288-11eb-8542-65af756c87c3.png?x-oss-process=image/resize,w_560 2 = 0.75

Confidence level = 0.90

α = 1 - 0.90 = 0.1 , 1 - (α/2) = 0.95

So z = 1.645 ----( from z score table )

=(0.28-0.25)\pm 1.645*\sqrt{\frac{(0.28*0.72)}{180}+\frac{(0.25*0.75)}{250}}

= -0.041 and 0.101

90% confidence interval for P1-P2 is -0.041, 0.101

c)

11 = 100, 2 = 120 , //img.homeworklib.com/questions/28e1c2a0-b288-11eb-882f-0923a640395e.png?x-oss-process=image/resize,w_560 1= 0.47, //img.homeworklib.com/questions/292024b0-b288-11eb-ae58-bfa2d3f54989.png?x-oss-process=image/resize,w_560 2 = 0.59 , //img.homeworklib.com/questions/295c52e0-b288-11eb-82c8-7d9cde331299.png?x-oss-process=image/resize,w_560 1 = 1 – //img.homeworklib.com/questions/29bb8e20-b288-11eb-9311-f7352a6b4900.png?x-oss-process=image/resize,w_560 1 = 0.53 , //img.homeworklib.com/questions/29f830c0-b288-11eb-953d-176cbb53a276.png?x-oss-process=image/resize,w_560 2 = 1 – //img.homeworklib.com/questions/2a350180-b288-11eb-8542-65af756c87c3.png?x-oss-process=image/resize,w_560 2 = 0.41

Confidence level = 0.90

α = 1 - 0.90 = 0.1 , 1 - (α/2) = 0.95

So z = 1.645 ----( from z score table )

=(0.47-0.59)\pm 1.645*\sqrt{\frac{(0.47*0.53)}{100}+\frac{(0.59*0.41)}{120}}

= -0.23 and -0.01

90% confidence interval for P1-P2 is -0.23 , -0.01